OpenStudy (anonymous):

What is the magnitude and direction of GH if G(1, -6) and H(8, 3)?

OpenStudy (anonymous):

A) magnitude: 11.4 units: direction: 37.87 B) magnitude: 11.4 units; direction: 52.13° C) magnitude: 7.62 units; direction 37.87° D) magnitude: 7.62 units: direction: 52.13°

OpenStudy (lgbasallote):

first what is y2 - y1?

OpenStudy (anonymous):

9?

OpenStudy (lgbasallote):

correct...what is x2 - x1?

OpenStudy (anonymous):

7?

OpenStudy (lgbasallote):

good correct now use the formula \[magnitude = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] got it?

OpenStudy (anonymous):

11.401754 ?

OpenStudy (lgbasallote):

yup orjust simply 11.4 units

OpenStudy (lgbasallote):

now to find theta \[\theta = \tan^{-1} (\frac{y_2 - y_1}{x_2 - x_1})\]

OpenStudy (lgbasallote):

so \[\theta = \tan^{-1} (\frac97)\] got that?

OpenStudy (anonymous):

idk how to do that !

OpenStudy (lgbasallote):

you use a calculator :S

OpenStudy (lgbasallote):

there's a button there that looks like \(\tan^{-1}\)

OpenStudy (lgbasallote):

usually it's shift + tan do you see it?

OpenStudy (lgbasallote):

i have to go now

OpenStudy (lgbasallote):

i hope you figure it out :)

OpenStudy (anonymous):

its B . thank you (:

OpenStudy (anonymous):

sorry i cudn't solve it before

OpenStudy (anonymous):

i mean sooner